# The vast world of quantum advantage

### v1

@@ II.4 Verifiability
Example 11 (Classical verification of quantum computation Mahadev (2018)).
@@ II.5 Usefulness
The emphasis on practical value for users who don’t care about the underlying physics is crucial. Many early quantum algorithms focused on problems specifically constructed to showcase quantum advantages, like random circuit sampling Arute et al. (2019) or boson sampling Aaronson and Arkhipov (2011). While these demonstrate fascinating quantum-classical separations, their practical utility remains limited. In contrast, truly useful quantum advantages should address problems we genuinely want to solve, regardless of how we solve them.
@@
Quantum simulation of strongly correlated quantum materials exemplifies a potentially useful quantum advantage. Understanding and predicting the properties of complex materials like high-temperature superconductors, exotic quantum magnets, or catalyst molecules has immense scientific and technological value, independent of the methods used. Classical computers face fundamental challenges in simulating these systems due to the intricate quantum correlations present in many-body systems, while quantum computers can naturally represent and evolve these quantum states. However, the landscape of simulation methods continues to evolve rapidly. Classical algorithms, particularly those leveraging machine learning and tensor networks Orús (2014), have shown remarkable progress in describing previously intractable quantum systems. It remains challenging to predict whether future classical technological breakthroughs will enable efficient simulation of the quantum materials most relevant to practical applications.
@@ III Realms of quantum advantage
The vast world of quantum advantage extends far beyond computational speedups and presents a rich tapestry of opportunities where quantum resources enable capabilities beyond those achievable with classical resources alone. We can organize these quantum advantages into distinct realms, each characterized by fundamentally different sources of quantum superiority over classical approaches. These realms differ not only in their applications but in their fundamental nature. Some rely on unproven complexity assumptions, while others are guaranteed by the laws of physics themselves.
@@ III.1 Computational advantages
At the foundation of the computational realm lies Shor’s algorithm Shor (1997), which exemplifies genuine quantum advantage through the hidden subgroup structure that quantum computers can exploit efficiently. As discussed in earlier sections, it provides compelling evidence for quantum advantage that satisfies our keystone criteria. The algorithm’s threat to cryptographic infrastructure worldwide demonstrates how quantum advantage can have profound real-world implications. Even decades after its discovery, no efficient classical algorithm has emerged to challenge its speedup.
@@
Quantum simulation is a particularly natural computational application, as the task of simulating quantum many-body systems inherently involves quantum mechanical effects that classical computers struggle to capture efficiently Feynman (1982); Lloyd (1996). The intuition is compelling: quantum systems should naturally simulate quantum physics. Yet this intuitive advantage faces relentless competition from advancing classical methods. Sophisticated tensor network algorithms Orús (2014), machine learning approaches Carleo and Troyer (2017), and quantum-inspired classical methods Lee et al. (2023) continue to push the boundaries of classical simulation. While quantum simulators have the potential to achieve advantages for specific problems, the landscape remains dynamic. What appears quantum-advantageous today may be challenged by classical innovations tomorrow.
@@
A subtle challenge to computational quantum advantages emerges with the recognition of the power of data Huang et al. (2021b). When quantum experiments generate data that cannot be efficiently produced classically, even classical machine learning algorithms trained on such data can solve computational problems beyond the reach of classical algorithms operating in isolation Huang et al. (2022a); Lewis et al. (2024). This reveals a deeper vulnerability: as we accumulate more data about our quantum universe through quantum simulations, quantum sensors, and quantum experiments, classical machine learning algorithms, by exploiting that data, might learn to approximate quantum phenomena without requiring additional quantum computation. However, assuming $\mathsf{BPP}\neq\mathsf{BQP}$, classical algorithms cannot universally simulate all quantum experiments. This suggests that important applications can be realized by quantum computing and classical AI working together; for example, quantum computers might collect data in previously unexplored physical regimes, empowering classical AI to predict properties of exotic quantum systems that have not yet been observed.
The realm of computational advantages thus presents a nuanced picture: while the quantum advantage exhibited by foundational algorithms like Shor’s remains well grounded, the broader landscape is characterized by ongoing competition between quantum and classical computers. Quantum computational advantages require constant vigilance and diligent verification as the classical state-of-the-art continues to advance.
@@ III.2 Learning/sensing advantages
Quantum sensing harnesses intrinsically quantum objects, such as atoms, ions, molecules, or photons, to measure physical quantities with enhanced performance. This quantum enhancement requires sophisticated quantum states, exemplified by GHZ states for magnetic field sensing Giovannetti et al. (2004, 2006). As described in Section II.3, achieving this quantum advantage faces a fundamental adversary: noise. The Hamiltonian-not-in-Lindbladian-Span (HNLS) criterion reveals that a quantum enhancement survives in the presence of noise only when the target signal and noise processes can be distinguished at the operator level Zhou et al. (2018). In generic scenarios, this condition fails, rendering the asymptotic quantum sensing advantage fundamentally unattainable even with quantum error correction Demkowicz-Dobrzański et al. (2012). This stands in stark contrast to quantum computation, where universal fault tolerance provides a clear path to noise robustness.
@@
Despite theoretical limitations, quantum sensing has already achieved remarkable real-world successes. While finding examples where entanglement enhances sensing capabilities in a real-world setting has been famously elusive, a notable exception is LIGO’s detection of gravitational waves, employing squeezed light to reduce quantum noise below the standard quantum limit Collaboration (2011); Abbott et al. (2016); Schnabel (2017); Ganapathy et al. (2023). In contrast, advantages exploiting the innate sensitivity of quantum sensors or ensembles thereof have been far more common. For example, nitrogen-vacancy centers in diamond enable quantum-enhanced magnetometry with nanoscale spatial resolution Taylor et al. (2008); Lovchinsky et al. (2016) or under extreme temperature or pressure conditions Bhattacharyya et al. (2024); Wang et al. (2024), while quantum-enhanced atomic clocks push temporal precision to unprecedented levels Ludlow et al. (2015); Bothwell et al. (2022); Oelker et al. (2022); Shaniv et al. (2024); Young et al. (2024). In sensing applications, even constant factor improvements over existing technology can provide transformative practical utility. Unlike computational tasks, where performance can often be enhanced by distributing workloads across multiple processors, sensing tasks frequently face fundamental constraints imposed by their physical infrastructure. For instance, achieving nanoscale spatial resolution in magnetometry requires the sensor to be physically positioned at the target location, while gravitational wave detection demands interferometers with enormous 4-kilometer-long arms. These physical constraints make it prohibitively expensive or physically impossible to simply deploy additional sensors to enhance performance. Quantum enhancements thus offer a fundamentally different value proposition than computational speedups.
@@
Recent developments have found that the most significant advantages emerge when we consider quantum machines that can see with quantum sensors, think with quantum computers, and remember with quantum memory Huang et al. (2021a); Aharonov et al. (2022); Chen et al. (2022); Huang et al. (2022a); Chen et al. (2024b); Oh et al. (2024); Allen et al. (2025); Liu et al. (2025a). Such quantum learning agents signify a paradigm shift beyond traditional quantum sensing approaches by enabling coherent processing of quantum information across multiple experimental runs.
@@
When applied to sensing classical oscillating fields such as electromagnetic signals or gravitational waves, quantum agents could achieve beyond-quadratic polynomial speedups. While quantum algorithms like Grover search Grover (1996) were not traditionally considered relevant to sensing applications, recent work Allen et al. (2025) demonstrates how quantum computers can be integrated with quantum sensors in a framework termed quantum-computing-enhanced sensing to learn properties of classical fields substantially faster than quantum sensors controlled by classical computers alone.
@@
While the aforementioned advantages are polynomial, one may wonder whether exponential quantum advantages exist in the learning and sensing realm. The answer is proven to be affirmative when quantum learning agents are applied to study many-body physical systems and dynamics Huang et al. (2021a); Aharonov et al. (2022); Chen et al. (2022); Huang et al. (2022a); Chen et al. (2024b); Oh et al. (2024); Liu et al. (2025a). These quantum agents can learn from exponentially fewer experiments than classical approaches for tasks including predicting physical properties and learning approximate models of quantum dynamics. The exponential improvement emerges from the ability to preserve quantum correlations across multiple experiments that classical approaches must necessarily destroy through intermediate measurements, enabling quantum learning agents to extract exponentially more information per experiment than their classical counterparts.
Rather than relying on unproven complexity assumptions, quantum advantages in learning/sensing realm often emerge from the quantum nature of physical reality itself. Looking ahead, the most transformative opportunities may emerge from quantum learning agents that seamlessly integrate quantum sensors and quantum computers. Some of these advantages have been proven to persist under realistic noise Huang et al. (2022a, b); Chen et al. (2023); Liu et al. (2025a), but the general theory of fault-tolerant quantum agents remains to be developed.
@@ III.3 Cryptographic/communication/strategic game advantages
Quantum mechanics enables parties to establish secure communication channels and verify information authenticity with guarantees that transcend computational security. Quantum key distribution exemplifies this revolutionary capability: the BB84 protocol Bennett and Brassard (1984) enables Alice and Bob to establish a shared cryptographic key via quantum communication with security guaranteed by the laws of physics. Any eavesdropper necessarily disturbs the quantum transmission, revealing their presence through detectable changes in the quantum states. This represents a paradigmatic shift from classical cryptographic protocols, where security relies on the assumption that computations that could break the protocol in principle are too difficult to execute in practice. Such computational assumptions could potentially be undermined by future algorithmic breakthroughs or advances in computational power. Certified randomness generation Colbeck and Renner (2012); Liu et al. (2025b) exploits these principles to create truly random bits that have not previously been generated prior to the protocol, whose unpredictability is guaranteed by quantum mechanics. Unlike classical pseudorandom number generators that are fundamentally deterministic, quantum randomness provides bits that cannot be spoofed by a secret seed or other backdoor. Furthermore, the randomness is guaranteed even if the user does not trust the inner workings of the quantum device used in the protocol. Such device-independent randomness certification has profound implications for cryptographic applications requiring secure random bits.
@@
The no-cloning theorem enables the creation of objects with inherent copy-protection, opening new classes of applications impossible in classical physics. Quantum money schemes Wiesner (1983) aim to create unforgeable currency by using quantum states to encode serial numbers that can be efficiently verified but cannot be perfectly duplicated. Certified deletion protocols Broadbent and Islam (2020) enable Alice to send quantum-encrypted data to Bob, who can later provide cryptographic proof that he has permanently and verifiably destroyed the information. This capability is impossible with classical physics, where information can always be copied before deletion. Copy-protected software Aaronson (2009) and one-time programs Goldwasser et al. (2008) extend these principles to software distribution, enabling programs that cannot be duplicated while preserving their computational functionality. These protocols may revolutionize intellectual property protection, data privacy, regulatory compliance, and rights management by making such violations physically impossible rather than merely legally prohibited.
@@
Quantum entanglement transcends classical communication limitations through nonlocal correlations that exhibit behaviors impossible in any local classical theory. As explored in Puzzle I, entangled quantum systems demonstrate correlations that cannot be reproduced by any classical mechanism. In strategic scenarios requiring spatially separated parties to act simultaneously in a coordinated manner, quantum protocols could provide fundamental advantages over classical protocols. For example, Ding and Jiang (2024) has illustrated how entangled quantum systems shared between spatially separated stock exchanges can provide coordination advantages in some high-frequency trading applications, where small time improvements can translate into significant competitive value.
@@ III.4 Space advantages
As demonstrated in Puzzle I, quantum states can serve as exponentially compressed representations of classical data for specific computational tasks. While Holevo’s bound Holevo (1973) forbids the extraction of an exponential number of classical bits from a small quantum memory, certain computational tasks can be performed on the compressed quantum representation without full data reconstruction. This quantum compression becomes particularly powerful in streaming scenarios where massive data streams must be processed with limited memory. Classical algorithms must make irreversible decisions about which information to retain as each data element arrives. In contrast, quantum algorithms can maintain superpositions over many possible configurations simultaneously, effectively postponing these irreversible decisions until more information becomes available.
@@
Recent work on directed maximum cut problems Kallaugher et al. (2024) demonstrates how quantum algorithms can solve natural graph problems using exponentially less memory than classical approaches when the input size vastly exceeds memory capacity. The vector-in-subspace problem studied by Raz Raz (1999) provides another example: determining which of two subspaces contains a given vector requires exponential classical communication but only linear quantum communication. However, space advantages face important limitations. They typically apply only when input size significantly exceeds memory capacity; furthermore, inputting a large amount of data into a quantum system can be costly, and the computational task performed must be compatible with Holevo’s bound. Despite these constraints, space advantages represent a promising direction as data generation accelerates.
@@ IV.3 Unpredictable quantum advantages
To establish this limitation rigorously, we consider a concrete computational problem regarding whether a specified computational task admits a quantum advantage or not. This problem is inspired by the field of meta-complexity, the study of the complexity of computational problems which address questions about complexity theory itself Ren and Santhanam (2023); Simons Institute for the Theory of Computing (2023). Earlier, we discussed how Pauli propagation provides an efficient classical method for simulating most quantum circuits, though large families of circuits remain classically intractable. This naturally leads us to ask: given a specific quantum circuit, can we determine whether executing it on a quantum computer would outperform this classical simulation method?
@@
This unpredictability extends beyond computational problems to the frontiers of scientific discovery. Quantum technologies are already kindling new insights into the behavior of highly entangled quantum many-body systems; recent examples include quantum many-body scars and deep thermalization Bernien et al. (2017); Turner et al. (2018); Jia et al. (2023); Sierant and Schirò (2023). What marvelous phenomena are yet to be found? We cannot possibly know if our classical technologies are intrinsically incapable of predicting the phenomena. Thus a rich panoply of surprises may be unveiled as we probe the physical universe with ever more advanced quantum devices.

### v2

@@ II.4 Verifiability
Example 12 (Classical verification of quantum computation [18]).
@@ II.5 Usefulness
The emphasis on practical value for users who don’t care about the underlying physics is crucial. Many early quantum algorithms focused on problems specifically constructed to showcase quantum advantages, like random circuit sampling [42] or boson sampling [62]. While these demonstrate fascinating quantum-classical separations, their practical utility remains limited. In contrast, truly useful quantum advantages should address problems we genuinely want to solve, regardless of how we solve them.
@@
Quantum simulation of strongly correlated quantum materials exemplifies a potentially useful quantum advantage. Understanding and predicting the properties of complex materials like high-temperature superconductors, exotic quantum magnets, or catalyst molecules has immense scientific and technological value, independent of the methods used. Classical computers face fundamental challenges in simulating these systems due to the intricate quantum correlations present in many-body systems, while quantum computers can naturally represent and evolve these quantum states. However, the landscape of simulation methods continues to evolve rapidly. Classical algorithms, particularly those leveraging machine learning and tensor networks [63], have shown remarkable progress in describing previously intractable quantum systems. It remains challenging to predict whether future classical technological breakthroughs will enable efficient simulation of the quantum materials most relevant to practical applications.
@@ III Realms of quantum advantage
The vast world of quantum advantage extends far beyond computational speedups and presents a rich tapestry of opportunities where quantum resources enable capabilities beyond those achievable with classical resources alone. We can organize these quantum advantages into distinct realms, each characterized by fundamentally different sources of quantum superiority over classical approaches. These realms differ not only in their applications but in their fundamental nature. Some rely on unproven complexity assumptions, while others are guaranteed by the laws of physics themselves. For comprehensive surveys of quantum algorithms and protocols across these realms, we refer readers to existing resources [64, 65, 66, 67].
@@ III.1 Computational advantages
Oracle-based advantages.
The theoretical power of quantum computation was first established in oracular settings [68, 69], where an oracle for computing a function is given as a black box. Simon’s algorithm [69] provided the first exponential advantage and directly inspired the development of Shor’s algorithm. Other foundational results include Grover’s search [70] and quantum walks [71, 72, 73] exhibiting polynomial to exponential oracular advantages. While oracle-based advantages provide important conceptual foundations for understanding quantum computational power, they are not computational advantages in the usual sense. The black box assumption means that when the oracle’s internal structure is revealed, the quantum advantage could disappear as classical algorithms may find efficient shortcuts that exploit the revealed structure.
@@
Building on the theoretical insights from Simon’s algorithm, Shor’s algorithm [16] exemplifies genuine computational quantum advantage. It exploits the hidden subgroup structure that quantum computers can efficiently solve, providing a concrete path from oracle-based advantage to real-world impact. As discussed in earlier sections, it satisfies our keystone criteria and demonstrates how quantum advantage can have profound real-world implications through its threat to cryptographic infrastructure worldwide. Even decades after its discovery, no efficient classical algorithm has emerged to challenge its speedup.
@@
Beyond number-theoretic problems, quantum simulation represents a particularly natural computational application, as the task of simulating quantum many-body systems inherently involves quantum mechanical effects that classical computers struggle to capture efficiently [74, 75]. The intuition is compelling: quantum systems should naturally simulate quantum physics. Yet this intuitive advantage faces relentless competition from advancing classical methods. Sophisticated tensor network algorithms [63], machine learning approaches [76], and quantum-inspired classical methods [77] continue to push the boundaries of classical simulation. While quantum simulators have the potential to achieve advantages for specific problems, the landscape remains dynamic. What appears quantum-advantageous today may be challenged by classical innovations tomorrow.
@@
A subtle challenge to computational quantum advantages emerges with the recognition of the power of data [78]. When quantum experiments generate data that cannot be efficiently produced classically, even classical machine learning algorithms trained on such data can solve computational problems beyond the reach of classical algorithms operating in isolation [79, 80]. This reveals a deeper vulnerability: as we accumulate more data about our quantum universe through quantum simulations, quantum sensors, and quantum experiments, classical machine learning algorithms, by exploiting that data, might learn to approximate quantum phenomena without requiring additional quantum computation. However, assuming $\mathsf{BPP}\neq\mathsf{BQP}$, classical algorithms cannot universally simulate all quantum experiments. This suggests that important applications can be realized by quantum computing and classical AI working together; for example, quantum computers might collect data in previously unexplored physical regimes, empowering classical AI to predict properties of exotic quantum systems that have not yet been observed.
The realm of computational advantages presents a nuanced picture: while the quantum advantage exhibited by foundational algorithms like Shor’s remains well grounded, the broader landscape is characterized by ongoing competition between quantum and classical computers. Quantum computational advantages require constant vigilance as the classical state-of-the-art continues to advance.
@@ III.2 Learning/sensing advantages
Quantum sensing harnesses intrinsically quantum objects, such as atoms, ions, molecules, or photons, to measure physical quantities with enhanced performance. This quantum enhancement requires sophisticated quantum states, exemplified by GHZ states for magnetic field sensing [81, 82]. As described in Section II.3, achieving this quantum advantage faces a fundamental adversary: noise. The Hamiltonian-not-in-Lindbladian-Span (HNLS) criterion reveals that a quantum enhancement survives in the presence of noise only when the target signal and noise processes can be distinguished at the operator level [52]. In generic scenarios, this condition fails, rendering the asymptotic quantum sensing advantage fundamentally unattainable even with quantum error correction [83]. This stands in stark contrast to quantum computation, where universal fault tolerance provides a clear path to noise robustness.
@@
Despite theoretical limitations, quantum sensing has already achieved remarkable real-world successes. While finding examples where entanglement enhances sensing capabilities in a real-world setting has been famously elusive, a notable exception is LIGO’s detection of gravitational waves, employing squeezed light to reduce quantum noise below the standard quantum limit [84, 85, 86, 87]. In contrast, advantages exploiting the innate sensitivity of quantum sensors or ensembles thereof have been far more common. For example, nitrogen-vacancy centers in diamond enable quantum-enhanced magnetometry with nanoscale spatial resolution [88, 89] or under extreme temperature or pressure conditions [90, 91], while quantum-enhanced atomic clocks push temporal precision to unprecedented levels [92, 93, 94, 95]. In sensing applications, even constant factor improvements over existing technology can provide transformative practical utility. Unlike computational tasks, where performance can often be enhanced by distributing workloads across multiple processors, sensing tasks frequently face fundamental constraints imposed by their physical infrastructure. For instance, achieving nanoscale spatial resolution in magnetometry requires the sensor to be physically positioned at the target location, while gravitational wave detection demands interferometers with enormous 4-kilometer-long arms. These physical constraints make it prohibitively expensive or physically impossible to simply deploy additional sensors to enhance performance. Quantum enhancements thus offer a fundamentally different value proposition than computational speedups.
@@
Recent developments have found that the most significant advantages emerge when we consider quantum machines that can see with quantum sensors, think with quantum computers, and remember with quantum memory [56, 96, 57, 58, 97, 98, 99, 61]. Such quantum learning agents signify a paradigm shift beyond traditional quantum sensing approaches by enabling coherent processing of quantum information across multiple experimental runs.
@@
When applied to sensing classical oscillating fields such as electromagnetic signals or gravitational waves, quantum agents could achieve beyond-quadratic polynomial speedups. While quantum algorithms like Grover search [70] were not traditionally considered relevant to sensing applications, recent work [99] demonstrates how quantum computers can be integrated with quantum sensors in a framework termed quantum-computing-enhanced sensing to learn properties of classical fields substantially faster than quantum sensors controlled by classical computers alone.
@@
While the aforementioned advantages are polynomial, one may wonder whether exponential quantum advantages exist in the learning and sensing realm. The answer is proven to be affirmative when quantum learning agents are applied to study many-body physical systems and dynamics [56, 96, 57, 58, 97, 98, 61]. These quantum agents can learn from exponentially fewer experiments than classical approaches for tasks including predicting physical properties and learning approximate models of quantum dynamics. The exponential improvement emerges from the ability to preserve quantum correlations across multiple experiments that classical approaches must necessarily destroy through intermediate measurements, enabling quantum learning agents to extract exponentially more information per experiment than their classical counterparts.
Rather than relying on unproven complexity assumptions, quantum advantages in learning/sensing realm often emerge from the quantum nature of physical reality itself. Looking ahead, the most transformative opportunities may emerge from quantum learning agents that seamlessly integrate quantum sensors and quantum computers. Some of these advantages have been proven to persist under realistic noise [58, 59, 60, 61], but the general theory of fault-tolerant quantum agents remains to be developed.
@@ III.3 Cryptographic/communication/strategic game advantages
Quantum mechanics enables parties to establish secure communication channels and verify information authenticity with guarantees that transcend computational security. Quantum key distribution exemplifies this revolutionary capability: the BB84 protocol [100] enables Alice and Bob to establish a shared cryptographic key via quantum communication with security guaranteed by the laws of physics. Any eavesdropper necessarily disturbs the quantum transmission, revealing their presence through detectable changes in the quantum states. This represents a paradigmatic shift from classical cryptographic protocols, where security relies on the assumption that computations that could break the protocol in principle are too difficult to execute in practice. Such computational assumptions could potentially be undermined by future algorithmic breakthroughs or advances in computational power. Certified randomness generation [101, 102] exploits these principles to create truly random bits that have not previously been generated prior to the protocol, whose unpredictability is guaranteed by quantum mechanics. Unlike classical pseudorandom number generators that are fundamentally deterministic, quantum randomness provides bits that cannot be spoofed by a secret seed or other backdoor. Furthermore, the randomness is guaranteed even if the user does not trust the inner workings of the quantum device used in the protocol. Such device-independent randomness certification has profound implications for cryptographic applications requiring secure random bits.
@@
The no-cloning theorem enables the creation of objects with inherent copy-protection, opening new classes of applications impossible in classical physics. Quantum money schemes [103] aim to create unforgeable currency by using quantum states to encode serial numbers that can be efficiently verified but cannot be perfectly duplicated. Certified deletion protocols [104] enable Alice to send quantum-encrypted data to Bob, who can later provide cryptographic proof that he has permanently and verifiably destroyed the information. This capability is impossible with classical physics, where information can always be copied before deletion. Copy-protected software [105] and one-time programs [106] extend these principles to software distribution, enabling programs that cannot be duplicated while preserving their computational functionality. These protocols may revolutionize intellectual property protection, data privacy, regulatory compliance, and rights management by making such violations physically impossible rather than merely legally prohibited.
@@
Quantum entanglement transcends classical communication limitations through nonlocal correlations that exhibit behaviors impossible in any local classical theory. As explored in Puzzle I, entangled quantum systems demonstrate correlations that cannot be reproduced by any classical mechanism. In strategic scenarios requiring spatially separated parties to act simultaneously in a coordinated manner, quantum protocols could provide fundamental advantages over classical protocols. For example, [107] has illustrated how entangled quantum systems shared between spatially separated stock exchanges can provide coordination advantages in some high-frequency trading applications, where small time improvements can translate into significant competitive value.
@@ III.4 Space advantages
As demonstrated in Puzzle I, quantum states can serve as exponentially compressed representations of classical data for specific computational tasks. While Holevo’s bound [7] forbids the extraction of an exponential number of classical bits from a small quantum memory, certain computational tasks can be performed on the compressed quantum representation without full data reconstruction. This quantum compression becomes particularly powerful in streaming scenarios where massive data streams must be processed with limited memory. Classical algorithms must make irreversible decisions about which information to retain as each data element arrives. In contrast, quantum algorithms can maintain superpositions over many possible configurations simultaneously, effectively postponing these irreversible decisions until more information becomes available.
@@
Recent work on directed maximum cut problems [108] demonstrates how quantum algorithms can solve natural graph problems using exponentially less memory than classical approaches when the input size vastly exceeds memory capacity. The vector-in-subspace problem studied by Raz [8] provides another example: determining which of two subspaces contains a given vector requires exponential classical communication but only linear quantum communication. However, space advantages face important limitations. They typically apply only when input size significantly exceeds memory capacity; furthermore, inputting a large amount of data into a quantum system can be costly, and the computational task performed must be compatible with Holevo’s bound. Despite these constraints, space advantages represent a promising direction as data generation accelerates.
@@ IV.3 Unpredictable quantum advantages
To establish this limitation rigorously, we consider a concrete computational problem regarding whether a specified computational task admits a quantum advantage or not. This problem is inspired by the field of meta-complexity, the study of the complexity of computational problems which address questions about complexity theory itself [109, 110]. Earlier, we discussed how Pauli propagation provides an efficient classical method for simulating most quantum circuits, though large families of circuits remain classically intractable. This naturally leads us to ask: given a specific quantum circuit, can we determine whether executing it on a quantum computer would outperform this classical simulation method?
@@
This unpredictability extends beyond computational problems to the frontiers of scientific discovery. Quantum technologies are already kindling new insights into the behavior of highly entangled quantum many-body systems; recent examples include quantum many-body scars and deep thermalization [111, 112, 113, 114]. What marvelous phenomena are yet to be found? We cannot possibly know if our classical technologies are intrinsically incapable of predicting the phenomena. Thus a rich panoply of surprises may be unveiled as we probe the physical universe with ever more advanced quantum devices.